OpenStudy (anonymous):

Which expression is equivalent to 2(m − 4) + 1?

OpenStudy (anonymous):

2m-4

OpenStudy (anonymous):

here are the expressions 2m - 7 2m - 3 2m + 9 2m + 10

OpenStudy (anonymous):

it has to be one of those

OpenStudy (anonymous):

ok hold on

OpenStudy (♪chibiterasu):

First distribute the 2 into the parenthesis.

OpenStudy (♪chibiterasu):

 By multiplying.

OpenStudy (♪chibiterasu):

Can you do that?

OpenStudy (anonymous):

so its 2m + 9

OpenStudy (♪chibiterasu):

Nope! When adding a positive to a negative, the negative will decrease in size.

OpenStudy (♪chibiterasu):

For example: -4 + 2 = -2 -3 + 1 = -2

OpenStudy (anonymous):

i got 2m-7

OpenStudy (♪chibiterasu):

Yeah, that.

OpenStudy (anonymous):

yep got that too thank you guys

OpenStudy (♪chibiterasu):

Np.

OpenStudy (anonymous):

no probz lil buddy